OpenStudy (anonymous):

cos^2x + 2 cos x + 1 = 0

OpenStudy (anonymous):

The first step would be to factor this, just like we factor any other polynomial. We're looking for two values that multiply to give 1 but add to give 2. These would be 1 and 1: \[(\cos x + 1)(\cos x + 1)=0\] So, since we have one factor, we set it equal to 0: \[\cos x + 1 = 0\] Subtract 1 from both sides \[\cos x = -1\] Now we solve. Cosine is equal to -1 at 180 degrees or pi radians.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

You're welcome, glad I could help.